# Towards Global AI Inclusivity: A Large-Scale Multilingual Terminology Dataset

### v1

@@ 4.2 Task 2: Is Our Dataset Translation Better than 60-60?
The 60-60 initiative dataset also focuses on AI terminology translations and includes the five languages we considered. This overlap allows us to intersect the English terms in our dataset with those in the 60-60 evaluation dataset and compare their respective translations. We retrieved the translations from both datasets and conducted a comparative assessment through a crowdsourced evaluation, same as the process detailed in Section 4.1.
@@
5 Dataset Coverage Assessments
@@
We first investigate the AI domains covered by the collected terms in GIST and their distributions. We use GPT-4o-mini (Hurst et al., 2024) to identify the specific AI domains for each term, following the taxonomy proposed by Ding et al. (2023), which clusters the research domains of AI scholars. Figure 8 shows the distribution of the top six AI domains in GIST, with terms most frequently categorized into statistics, mathematics, computer science (CS), natural language processing (NLP), data science (DS), and computer vision (CV). We further embed all AI terminologies in GIST using the all-MiniLM-L6-v2 model (Reimers and Gurevych, 2019) and apply Uniform Manifold Approximation and Projection (UMAP) (McInnes et al., 2018) for dimensionality reduction to a two-dimensional space. The visualization in Figure 9 reveals that terms in domains such as NLP, CV, statistics, and mathematics form distinct clusters, while terms in CS and DS are more dispersed.
@@
We claim that our awarded paper dataset provides essential coverage of the terms in our terminology dictionary, where many terms are actually repeatedly extracted multiple times from different awarded papers. We present the rarefaction curve in Figure 10, which visually demonstrates the relationship between the number of extracted terms and the dataset size.
To statistically validate our claim, we conducted a one-sided one-sample t-test to compare the mean coverage ratio of the terminology dictionary. Specifically, we tested whether the mean coverage ratio of randomly sampled subsets, constituting 60% of the terminology dictionary, is significantly above 80%. The null hypothesis ($H_{0}$) assumes that the mean coverage ratio of these subsets is less than or equal to 80%, while the alternative hypothesis ($H_{a}$) assumes that the mean coverage ratio exceeds 80%. Using 1,000 random samples, the analysis yielded a t-statistic of 64.78 and a p-value of 0, rejecting the null hypothesis. These results provide statistical evidence supporting the sufficiency of our dataset in covering the terminology dictionary.
@@ 6.1 Methods
We apply the word alignment approach introduced by Dou and Neubig (2021) to identify and substitute term translations within the output. We use the multilingual BERT base model (Devlin et al., 2018) to to tokenize both source and target sentences, and then processes these inputs through its hidden layers to produce contextualized embeddings. We determine alignments by computing dot-product similarities between source and target token embeddings. High-confidence alignments are then filtered using a threshold of 1e-4, and subword alignments are aggregated to generate word-level mappings. Lastly, we directly replace identified term translations with those from GIST.
@@
For constrained beam search, we use the Hugging Face implementation, which enforces the inclusion of term translations from GIST in the output. [^5] https://huggingface.co/blog/constrained-beam-search For token-level adjustment, we first identify AI terminologies and their corresponding expected translations from GIST. We then modify the output logits of these tokens by increasing them by factors of 10/7, 10/8, and 10/9 to prioritize their selection during decoding.
@@ 6.2 Experiment Setup
We conducted experiments on two evaluation sets: the 60-60 Evaluation Set and the AI Papers and Model Cards Evaluation Set. The first evaluation set is from the 60-60 initiative, providing terminology translations into five target languages. Based on our analysis in Section 4.2, which indicates that our dataset offers higher quality than the 60-60 set, we enhanced their labels by replacing AI terminologies in their translations with GPT-4o while maintaining grammatical correctness and ensuring no loss in translation quality. This process generated updated ground truth labels for the evaluation.
The second evaluation set was manually created by combining text from two sources: 50 held-out AI research papers and 50 model cards generated by (Liu et al., 2024). From this set, we randomly sampled 500 English text chunks from it for evaluation. To create ground truth labels for this set, we used Google Translate, a state-of-the-art machine translation model (Zhu et al., 2023; Santosa et al., 2024), to produce initial translations. We then prompted GPT-4o to refine these translations by updating AI terminology. This evaluation set was designed to explore the application of our AI terminology dictionary in two major domains: AI research papers and model cards.
@@
We evaluated the following models: gpt-4o-mini (Hurst et al., 2024), hf-seamless-m4t-large (Barrault et al., 2023), nllb-200-3.3B (Costa-jussà et al., 2022), aya-23-8B (Aryabumi et al., 2024), and aya-expanse-8B (Dang et al., 2024). While the prompting method applies to all models, the word alignment and decoding methods are not applicable to gpt-4o-mini, as it does not provide access to its model weights.
@@
We adopt BLEU (Papineni et al., 2002) and COMET metrics (Rei et al., 2020) for the quantitative evaluation for translation quality, following the methodology of Salesky et al. (2023). We utilize the wmt22-comet-da model (Rei et al., 2022).
@@ 6.3 Experiment Results
Figure 3: The interface demonstrates a paper example from (Tang et al., 2024) and introduces a new translation feature on the ACL Anthology website. Non-English speakers can click the “Translate” button at the bottom right of the webpage and select their preferred language for translation. The page dynamically displays two translations side by side: one using direct translation from a machine translation model (e.g., seamless in the figure) and the other enhanced with prompting-powered refinement applied to the left translation. Highlighted text indicates updated terminologies integrated from GIST, showcasing improved translation performance.
@@
The word alignment method demonstrates mixed performance, showing improvements for Chinese and Japanese translations but leading to declines for Arabic, French, and Russian. This discrepancy is likely due to linguistic differences: Chinese and Japanese exhibit minimal morphological changes, allowing straightforward substitution of terminology without disrupting surrounding syntax; In contrast, languages like Arabic often require agreement in gender, number, and syntactic roles, making noun replacement more challenging and prone to errors.
We also conducted a qualitative manual evaluation of the generation results produced by constrained beam search and logit adjustment methods. Unfortunately, both approaches exhibited extremely slow performance, running approximately 100 times slower than post-hoc methods. Furthermore, forcing specific word ids to appear in the output did not perform well, suffering from similar issues as the word alignment method in disrupting syntactical dependencies. During manual inspection, we observed that constrained beam search behaved similarly to the logit adjustment method with a large scaling factor: the generated sentences often failed to maintain proper grammar, and terminologies were frequently repeated multiple times within the same output. Conversely, with a small scaling factor in the logit adjustment method, the terms were often omitted entirely, demonstrating little to no effect on the output. These findings highlight the limitations in these traditional decoding approaches compared to the prompting-powered refinement method, as well as the importance of tailoring terminology integration approaches to the linguistic characteristics of target languages.
@@ 7 Website Demonstration
To facilitate real-world usage of our AI terminology dictionary, we built on the 60-60 initiative by modifying the ACL Anthology website layout and introducing a new terminology translation feature, [^6] https://acl6060.org/ as illustrated in Figure 3. The ACL Anthology website was chosen for this demonstration due to its extensive collection of AI-related research papers, making it an ideal platform to showcase the potential impact of our work. In this demonstration, translations of key terms are refined and standardized based on the terminology dictionary we developed, providing more accurate terminology translations. This enhancement represents a step toward improving access to AI knowledge for non-English speakers by offering a consistent and reliable translation system, and broadens the accessibility of AI research to a global audience.
@@ A.2 Terminology Collection Details
We integrate terms from the Wiki-Pedia Glossary of AI777en.wikipedia.org/wiki/Glossary_of_artificial_intelligence, which serves as a comprehensive starting point for artificial intelligence (AI)-related terminology, ensuring alignment with globally recognized concepts and definitions. To enhance multilingual coverage and domain relevance, we also include terms from several other specialized AI terminology initiatives across different languages:
Arabic AI Dictionary888ai.gov.ae/ar/ai-dictionary/: Published by the Arabic Government AI Office, this resource aims to elevate the status of the Arabic language in AI, standardize terminology, reduce linguistic ambiguity, and foster better integration of Arabic speakers into the global AI community. By providing accurate translations and clear definitions for English AI terms, this dictionary promotes knowledge dissemination and encourages collaboration within the Arabic-speaking AI ecosystem.
@@ Appendix C Additional Experiment Results
We present comprehensive evaluation results using BLEU (Papineni et al., 2002), ChrF (Popović, 2015), ChrF++ (Popović, 2017), and TER (Snover et al., 2006) for translation into the five target languages across both evaluation sets.
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 22.38 23.75 23.02
Bleu score std 24.43 25.89 24.92
Bleu 1gram precision mean 22.48 23.85 23.13
Bleu 1gram precision std 24.47 25.92 24.96
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 70.18 70.31 70.33
Bleu ref len std 35.62 35.60 35.84
Bleu pred len mean 70.08 70.24 70.17
Bleu pred len std 35.40 35.41 35.62
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 22.87 24.24 23.51
Chrf score std 24.56 26.02 25.04
Ter score mean 91.12 89.53 90.37
Ter score std 29.39 31.08 29.94
Chrfpp score mean 22.87 24.24 23.51
Chrfpp score std 24.56 26.02 25.04
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 23.13 24.28 23.10
Bleu score std 24.72 25.84 24.57
Bleu 1gram precision mean 23.19 24.35 23.18
Bleu 1gram precision std 24.72 25.85 24.59
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 69.56 69.79 69.84
Bleu ref len std 35.32 35.54 35.56
Bleu pred len mean 69.64 69.89 69.83
Bleu pred len std 35.11 35.31 35.30
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 23.67 24.83 23.64
Chrf score std 24.92 26.04 24.78
Ter score mean 90.27 88.94 90.29
Ter score std 29.63 31.00 29.45
Chrfpp score mean 23.67 24.83 23.64
Chrfpp score std 24.92 26.04 24.78
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 23.58 24.64 23.57
Bleu score std 26.63 27.85 26.40
Bleu 1gram precision mean 23.62 24.69 23.64
Bleu 1gram precision std 26.64 27.86 26.43
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 70.40 70.59 70.67
Bleu ref len std 35.88 35.95 36.14
Bleu pred len mean 70.68 70.85 70.88
Bleu pred len std 35.92 36.00 36.14
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.01 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 24.06 25.13 24.06
Chrf score std 26.76 27.98 26.55
Ter score mean 89.66 88.49 89.54
Ter score std 32.09 33.43 31.74
Chrfpp score mean 24.06 25.13 24.06
Chrfpp score std 26.76 27.98 26.55
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 19.98 20.52 19.77
Bleu score std 22.35 23.49 22.14
Bleu 1gram precision mean 20.03 20.57 19.82
Bleu 1gram precision std 22.37 23.50 22.15
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 69.72 69.93 69.91
Bleu ref len std 35.74 36.02 35.93
Bleu pred len mean 69.99 70.20 70.14
Bleu pred len std 35.69 35.97 35.85
Bleu bp mean 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.01 1.01 1.01
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 20.50 21.03 20.28
Chrf score std 22.54 23.65 22.34
Ter score mean 93.97 93.40 94.20
Ter score std 26.96 28.24 26.69
Chrfpp score mean 20.50 21.03 20.28
Chrfpp score std 22.54 23.65 22.34
@@
Direct Translation Prompting Refinement Word Alignment
Bleu score mean 20.11 21.34 20.29
Bleu score std 23.09 24.68 23.43
Bleu 1gram precision mean 20.20 21.43 20.38
Bleu 1gram precision std 23.14 24.73 23.49
Bleu 2gram precision mean 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00
Bleu ref len mean 71.54 71.47 71.70
Bleu ref len std 36.10 36.29 36.28
Bleu pred len mean 71.79 71.75 71.93
Bleu pred len std 36.24 36.43 36.39
Bleu bp mean 0.99 0.99 0.99
Bleu bp std 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02
Chrf score mean 20.62 21.85 20.80
Chrf score std 23.26 24.83 23.60
Ter score mean 93.61 92.27 93.37
Ter score std 27.75 29.58 28.12
Chrfpp score mean 20.62 21.85 20.80
Chrfpp score std 23.26 24.83 23.60

### v3

@@ 4.2 Task 2: Is Our Dataset Translation Better than 60-60?
The 60-60 initiative dataset also focuses on AI terminology translations and includes the five languages we consider. This overlap allows us to intersect the English terms in our dataset with those in the 60-60 evaluation dataset and compare their respective translations. We retrieve the translations from both datasets and conduct a comparative assessment through a crowdsourced evaluation, same as the process detailed in Section 4.1.
@@
5 Dataset Coverage Assessment
@@
We first investigate the AI domains covered by the collected terms in GIST and their distributions. We use GPT-4o-mini (Hurst et al., 2024) to identify the specific AI domains for each term, following the taxonomy proposed by Ding et al. (2023), which clusters the research domains of AI scholars. Figure 8 shows the distribution of the top six AI domains in GIST, with terms most frequently categorized into statistics, mathematics, computer science (CS), natural language processing (NLP), data science (DS), and computer vision (CV). We further embed all AI terminology in GIST using the all-MiniLM-L6-v2 model (Reimers and Gurevych, 2019) and apply Uniform Manifold Approximation and Projection (UMAP) (McInnes et al., 2018) for dimensionality reduction to a two-dimensional space. The visualization in Figure 9 reveals that terms in domains such as NLP, CV, statistics, and mathematics form distinct clusters, while terms in CS and DS are more dispersed.
@@
We surmise that our awarded paper dataset provides essential coverage of the terms in our terminology dictionary, where many terms are actually repeatedly extracted multiple times from different awarded papers. We present the rarefaction curve in Figure 10, which visually demonstrates the relationship between the number of extracted terms and the dataset size.
To statistically validate our claim, we conduct a one-sided one-sample t-test to compare the mean coverage ratio of the terminology dictionary. Specifically, we test whether the mean coverage ratio of randomly sampled subsets, constituting 60% of the terminology dictionary, is significantly above 80%. The null hypothesis ($H_{0}$) assumes that the mean coverage ratio of these subsets is less than or equal to 80%, while the alternative hypothesis ($H_{a}$) assumes that the mean coverage ratio exceeds 80%. Using 1,000 random samples, the analysis yields a t-statistic of 64.78 and a p-value of 0, rejecting the null hypothesis. These results provide statistical evidence supporting the sufficiency of our dataset in covering the terminology dictionary.
@@ 6.1 Methods
We apply the word alignment approach introduced by Dou and Neubig (2021) to identify and substitute term translations within the output. We use the multilingual BERT base model (Devlin et al., 2018) to tokenize both source and target sentences, and then processes these inputs through its hidden layers to produce contextualized embeddings. We determine alignments by computing dot-product similarities between source and target token embeddings. High-confidence alignments are then filtered using a threshold of 1e-4, and subword alignments are aggregated to generate word-level mappings. Lastly, we directly replace identified term translations with those from GIST.
@@
For constrained beam search, we use the Hugging Face implementation, which enforces the inclusion of term translations from GIST in the output. [^5] https://huggingface.co/blog/constrained-beam-search For token-level adjustment, we first identify AI terminology and their corresponding expected translations from GIST. We then modify the output logits of these tokens by increasing them by factors of 10/7, 10/8, and 10/9 to prioritize their selection during decoding.
@@ 6.2 Experiment Setup
We conduct experiments on two evaluation sets: the 60-60 Evaluation Set and the AI Papers and Model Cards Evaluation Set. The first evaluation set is from the 60-60 initiative, providing terminology translations into five target languages. Based on our analysis in Section 4.2, which indicates that our dataset offers higher quality than the 60-60 set, we enhance their labels by replacing AI terminology in their translations with GPT-4o while maintaining grammatical correctness and ensuring no loss in translation quality. This process generates updated ground truth labels for the evaluation.
The second evaluation set is manually created by combining text from two sources: 50 held-out AI research papers and 50 model cards generated by (Liu et al., 2024). From this set, we randomly sampled 500 English text chunks from it for evaluation. To create ground truth labels for this set, we use Google Translate, a state-of-the-art machine translation model (Zhu et al., 2023; Santosa et al., 2024), to produce initial translations. We then prompt GPT-4o to refine these translations by updating AI terminology. This evaluation set is designed to explore the application of our AI terminology dictionary in two major domains: AI research papers and model cards.
@@
We evaluate the following models: gpt-4o-mini (Hurst et al., 2024), hf-seamless-m4t-large (Barrault et al., 2023), nllb-200-3.3B (Costa-jussà et al., 2022), aya-23-8B (Aryabumi et al., 2024), and aya-expanse-8B (Dang et al., 2024). While the prompting method applies to all models, the word alignment and decoding methods are not applicable to gpt-4o-mini, as it does not provide access to its model weights.
@@
We adopt BLEU (Papineni et al., 2002) and COMET metrics (Rei et al., 2020) for the quantitative evaluation of translation quality, following the methodology of Salesky et al. (2023). We utilize the wmt22-comet-da model (Rei et al., 2022).
@@ 6.3 Experiment Results
Figure 3: The interface demonstrates a paper example from (Tang et al., 2024) and introduces a new translation feature on the ACL Anthology website. Non-English speakers can click the “Translate” button at the bottom right of the webpage and select their preferred language for translation. The page dynamically displays two translations side by side: one using direct translation from a machine translation model (e.g., seamless in the figure) and the other enhanced with prompting-powered refinement applied to the left translation. Highlighted text indicates updated terminology integrated from GIST, showcasing improved translation performance.
@@
The word alignment method demonstrates mixed performance, showing improvements for Chinese and Japanese translations but leading to declines for Arabic, French, and Russian. This discrepancy is due to linguistic differences: Chinese and Japanese exhibit minimal morphological changes, allowing straightforward substitution of terminology without disrupting surrounding syntax; In contrast, languages like Arabic often require agreement in gender, number, and syntactic roles, making noun replacement more challenging and prone to errors. To validate this claim, we conduct a post-hoc prompting after word alignment using GPT-4o-mini to render the translations morphologically coherent and correct. Details of the prompting process are provided in Appendix B. As shown in Appendix C, this improves translation scores to align with the performance of the prompting-powered method across all languages and models.
We also conduct a qualitative manual evaluation of the generation results produced by constrained beam search and logit adjustment methods. Both approaches exhibit extremely slow performance, running approximately 100 times slower than post-hoc methods. Furthermore, forcing specific word ids to appear in the output did not perform well, suffering from similar issues as the word alignment method in disrupting syntactical dependencies and morphological agreements. During manual inspection, we observe that constrained beam search behaves similarly to the logit adjustment method with a large scaling factor: the generated sentences often failed to maintain proper grammar, and terminological terms were frequently repeated multiple times within the same output. Conversely, with a small scaling factor in the logit adjustment method, the terms were often omitted entirely, demonstrating little to no effect on the output. These findings highlight the limitations in these traditional decoding approaches compared to the prompt-powered refinement method, as well as the importance of tailoring terminology integration approaches to the linguistic characteristics of target languages.
@@ 7 Website Demonstration
To facilitate real-world usage of our AI terminology dictionary, we built on the 60-60 initiative by modifying the ACL Anthology website layout and introducing a new terminology translation feature, [^6] https://acl6060.org/ as illustrated in Figure 3. The ACL Anthology website was chosen for this demonstration due to its extensive collection of AI-related research papers, making it an ideal platform to showcase the potential impact of our work. In this demonstration, translations of key terms are refined and standardized based on the terminology dictionary we developed, providing more accurate terminology translations.777 Our code for the website is available at https://github.com/ImanOu123/acl-anthology-mod. This enhancement represents a step toward improving access to AI knowledge for non-English speakers by offering a consistent and reliable translation system, and broadens the accessibility of AI research to a global audience.
@@ 8 Conclusion
Acknowledgments
We extend our gratitude to the Minister of State for Artificial Intelligence of the United Arab Emirates for making the Arabic AI terminology dictionary publicly available. We thank DataFranca for granting us access to their French AI terminology dictionary. We thank Jiqizhixin for sharing their Chinese AI terminology database. We also thank the Russian Association of Artificial Intelligence for providing access to the Russian AI terminology dictionary. For more details, please refer to Section A.2.
@@ A.2 Terminology Collection Details
We integrate terms from the WikiPedia Glossary of AI888en.wikipedia.org/wiki/Glossary_of_artificial_intelligence, which serves as a comprehensive starting point for artificial intelligence (AI)-related terminology, ensuring alignment with globally recognized concepts and definitions. To enhance multilingual coverage and domain relevance, we also include terms from several other specialized AI terminology initiatives across different languages:
Arabic AI Dictionary999ai.gov.ae/ar/ai-dictionary/: Published by the Arabic Government AI Office, this resource aims to elevate the status of the Arabic language in AI, standardize terminology, reduce linguistic ambiguity, and foster better integration of Arabic speakers into the global AI community. By providing accurate translations and clear definitions for English AI terms, this dictionary promotes knowledge dissemination and encourages collaboration within the Arabic-speaking AI ecosystem.
@@ B.1 Terminology Refinement via Prompting Details
B.2 Terminology Substitution by Word Alignment Details
To validate our analysis in Section 6.3 regarding the varying performance of the word alignment method across different languages, we perform a post-hoc prompting step after word alignment. This step ensures that the translations are morphologically coherent and accurate, and is used to compare with the original results. The prompt is depicted in Figure 12.
Figure 12: Prompts for refining the output of the word alignment method to ensure morphological coherence and accuracy.
@@ Appendix C Additional Experiment Results
We present comprehensive evaluation results using BLEU (Papineni et al., 2002), ChrF (Popović, 2015), ChrF++ (Popović, 2017), TER (Snover et al., 2006), and COMET (Rei et al., 2020) for translation into the five target languages across both evaluation sets from Table 10 to Table 61. We also present one good example and one bad example for the prompting-powered refinement and the word alignment method across all languages, as shown from Figure 13 to Figure 17.
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 22.38 23.75 23.02 23.58
Bleu score std 24.43 25.89 24.92 25.48
Bleu 1gram precision mean 22.48 23.85 23.13 23.67
Bleu 1gram precision std 24.47 25.92 24.96 25.51
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 70.18 70.31 70.33 70.22
Bleu ref len std 35.62 35.60 35.84 35.67
Bleu pred len mean 70.08 70.24 70.17 70.25
Bleu pred len std 35.40 35.41 35.62 35.57
Bleu bp mean 0.99 0.99 0.99 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 22.87 24.24 23.51 24.08
Chrf score std 24.56 26.02 25.04 25.61
Ter score mean 91.12 89.53 90.37 89.78
Ter score std 29.39 31.08 29.94 30.63
Chrfpp score mean 22.87 24.24 23.51 24.08
Chrfpp score std 24.56 26.02 25.04 25.61
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 23.13 24.28 23.10 25.48
Bleu score std 24.72 25.84 24.57 27.26
Bleu 1gram precision mean 23.19 24.35 23.18 25.53
Bleu 1gram precision std 24.72 25.85 24.59 27.27
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 69.56 69.79 69.84 69.72
Bleu ref len std 35.32 35.54 35.56 35.55
Bleu pred len mean 69.64 69.89 69.83 69.89
Bleu pred len std 35.11 35.31 35.30 35.37
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.00
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 23.67 24.83 23.64 26.03
Chrf score std 24.92 26.04 24.78 27.44
Ter score mean 90.27 88.94 90.29 87.50
Ter score std 29.63 31.00 29.45 32.62
Chrfpp score mean 23.67 24.83 23.64 26.03
Chrfpp score std 24.92 26.04 24.78 27.44
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 23.58 24.64 23.57 24.08
Bleu score std 26.63 27.85 26.40 27.16
Bleu 1gram precision mean 23.62 24.69 23.64 24.13
Bleu 1gram precision std 26.64 27.86 26.43 27.18
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 70.40 70.59 70.67 70.61
Bleu ref len std 35.88 35.95 36.14 36.11
Bleu pred len mean 70.68 70.85 70.88 70.87
Bleu pred len std 35.92 36.00 36.14 36.14
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.01 1.00 1.00 1.01
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 24.06 25.13 24.06 24.56
Chrf score std 26.76 27.98 26.55 27.28
Ter score mean 89.66 88.49 89.54 89.12
Ter score std 32.09 33.43 31.74 32.65
Chrfpp score mean 24.06 25.13 24.06 24.56
Chrfpp score std 26.76 27.98 26.55 27.28
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 19.98 20.52 19.77 20.56
Bleu score std 22.35 23.49 22.14 23.37
Bleu 1gram precision mean 20.03 20.57 19.82 20.61
Bleu 1gram precision std 22.37 23.50 22.15 23.40
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 69.72 69.93 69.91 69.91
Bleu ref len std 35.74 36.02 35.93 35.83
Bleu pred len mean 69.99 70.20 70.14 70.21
Bleu pred len std 35.69 35.97 35.85 35.79
Bleu bp mean 1.00 1.00 1.00 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.01 1.01 1.01 1.01
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 20.50 21.03 20.28 21.08
Chrf score std 22.54 23.65 22.34 23.54
Ter score mean 93.97 93.40 94.20 93.36
Ter score std 26.96 28.24 26.69 28.17
Chrfpp score mean 20.50 21.03 20.28 21.08
Chrfpp score std 22.54 23.65 22.34 23.54
@@
Direct Translation Prompting Refinement Word Alignment Morphologically Correct Word Alignment
Bleu score mean 20.11 21.34 20.29 21.36
Bleu score std 23.09 24.68 23.43 24.39
Bleu 1gram precision mean 20.20 21.43 20.38 21.43
Bleu 1gram precision std 23.14 24.73 23.49 24.44
Bleu 2gram precision mean 0.00 0.00 0.00 0.00
Bleu 2gram precision std 0.00 0.00 0.00 0.00
Bleu 3gram precision mean 0.00 0.00 0.00 0.00
Bleu 3gram precision std 0.00 0.00 0.00 0.00
Bleu 4gram precision mean 0.00 0.00 0.00 0.00
Bleu 4gram precision std 0.00 0.00 0.00 0.00
Bleu ref len mean 71.54 71.47 71.70 71.36
Bleu ref len std 36.10 36.29 36.28 36.33
Bleu pred len mean 71.79 71.75 71.93 71.68
Bleu pred len std 36.24 36.43 36.39 36.47
Bleu bp mean 0.99 0.99 0.99 1.00
Bleu bp std 0.01 0.01 0.01 0.01
Bleu ratio mean 1.00 1.00 1.00 1.01
Bleu ratio std 0.02 0.02 0.02 0.02
Chrf score mean 20.62 21.85 20.80 21.87
Chrf score std 23.26 24.83 23.60 24.54
Ter score mean 93.61 92.27 93.37 92.26
Ter score std 27.75 29.58 28.12 29.27
Chrfpp score mean 20.62 21.85 20.80 21.87
Chrfpp score std 23.26 24.83 23.60 24.54